# PANDA – Paired Anti-hate Narratives Dataset from Asia: Using an LLM-as-a-Judge to Create the First Chinese Counterspeech Dataset

### v1

@@ Abstract
Despite the global prevalence of Modern Standard Chinese language, counterspeech (CS) resources for Chinese remain virtually nonexistent. To address this gap in East Asian counterspeech research we introduce the a corpus of Modern Standard Mandarin counterspeech that focuses on combating hate speech in Mainland China. This paper proposes a novel approach of generating CS by using an LLM-as-a-Judge, simulated annealing, LLMs zero-shot CN generation and a round-robin algorithm. This is followed by manual verification for quality and contextual relevance. This paper details the methodology for creating effective counterspeech in Chinese and other non-Eurocentric languages, including unique cultural patterns of which groups are maligned and linguistic patterns in what kinds of discourse markers are programmatically marked as hate speech (HS). Analysis of the generated corpora, we provide strong evidence for the lack of open-source, properly labeled Chinese hate speech data and the limitations of using an LLM-as-Judge to score possible answers in Chinese. Moreover, the present corpus serves as the first East Asian language based CS corpus and provides an essential resource for future research on counterspeech generation and evaluation.
Warning: The below text contains vulgar and oftentimes offensive speech. Any counter-speech or hate speech is used for exemplary purposes and doesn’t necessarily reflect the views of any researcher involved.
@@ 1 Introduction
Hate speech (HS) is typically characterized as any form of communication that demeans a specific group of people based on attributes like race, ethnicity, gender, sexual orientation, or religion (de Gibert et al., 2018). While HS may constitute a small proportion of social media content, its impact is significant, affecting nearly one-third of the population (Vidgen et al., 2019). The proliferation of hate speech on social media platforms has become a significant societal concern. While traditional approaches to mitigating HS have focused on content removal and moderation, these methods often raise concerns about freedom of speech. In response, counterspeech (CS) has emerged as a promising alternative strategy to combat HS while preserving free expression (Poudhar et al., 2024).
@@ 2.1 Hate Speech and Counterspeech
Counterspeech has gained traction as an alternative to content removal. Studies have demonstrated the efficacy of CS in enhancing online discourse quality and reducing the prevalence and impact of hateful behavior (Buerger, 2021). However, it’s important to note that the effectiveness of CS can vary significantly depending on the context and specific strategies employed. For example, the quantity of training data available to train an LLM on a specific language will predict the robustness of it’s generative function.
@@ 2.5 Current Evaluation Metrics
Datasets Open Source Total Instances HS/Offensive Speech Non-HS
@@ 3 Methodology
Figure 1: Data Processing Pipeline for Creating the Chinese Counterspeech Corpus. A1 through An refer to $n$ annotators that participated in this project.
@@ 3.1 Goals/Requirements
  • Creation of the First East Asian HS-CN Dataset. During our review of existing datasets, we identified significant gaps in Chinese counter-speech (CS) resources. Although datasets like Deng et al. (2022) and Zhou et al. (2022) include instances labeled as ’anti-bias’, their scope and definitions do not align with the specific focus of CS research. These datasets adopt a broader concept of ’anti-bias’, encompassing content that promotes fairness and addresses various forms of offensive language rather than specifically targeting hate speech. Our work addresses this gap by creating a dataset that exclusively targets hate speech and counter speech, providing a more focused resource for CS research.
@@
  • Freely Usable. All hate speech (HS) data collected for our dataset originate from open-source repositories. Additionally, we have released our model and the generated data under a permissive GPL license. This ensures that the generated and annotated data can be freely utilized in both commercial and non-commercial projects, promoting wider accessibility and application in various research and practical initiatives.
@@ 3.2 HS Sources
Two corpora ((Lu et al., 2023) and (Wang et al., 2022)) were later removed from the dataset due to restrictive licensing from them. One dataset (Zhou et al. (2022)) was not included due to a large majority of the potential HS text being included in the CHSD Rao et al. (2023).
What was left were 3 open-source datasets. The COLDataset contains over 30,000 instances that are labeled either safe or offensive and, further, contains fine grained labels for each category Deng et al. (2022). The dataset was chosen due the fact that, under a cursory look, many, but not all, of the statements labeled offensive were in-fact hate speech. The second dataset used was ‘SexComment.csv’ from SWSR. This file focuses on finding and labeling sexist comments and also contains subcategories for the type of comment and whether it is targeted at an individual or a group Jiang et al. (2022). We decided to include this dataset to increase representation of sexist hate-speech in the database. The last dataset included was from CHSD which is actually a pre-processed dataset of HS that comes from COLD, CDIAL, and SWSR Rao et al. (2023). We decided to use this corpus instead of CDIAL directly as it cursory looked to contain more relevant entries, but we decided to keep using SWSR and the COLDataset as the authors believed that there were still hate speech sentences from those sources that were excluded from CHSD.
@@ 3.3 Filtering of Data
The initial three corpora included entries that were labeled non-hate speech. In order to avoid the the unnecessary computational cost of attempting to generate CS for non-HS sentences, we initially used some commands to filter out any rows that aren’t considered HS by the corpora. For CSHD, we removed any rows where ‘label’ equals ‘0.’ Likewise, for the COLD dataset, we kept only the data that had a label of ‘1’ and a sub label of ‘1’ (attacking individuals) or ‘2’ (attacking groups). For SWSR, we keep only the instances that had a label of ‘1’ (sexist) and all sub-categories except for ‘MA’ (micro aggressions) as we believed that it was harder to determine which answers counted as a hate-speech.
@@
We then use the scores given by the the LLM and the text length to optimize over the set of possible subsets of hate-speech. As we wanted to have a subset that balances between high average hate score and a high average text length, we choose the metric of $log(AverageHateScore)*log(AverageTextLength)*NumInstances$. We limited the range from 500 to 3000 so that we would have a subset of answers that is large enough. As can be seen from Figure 2, we found that including strings that had a string length of at-least 53 characters and a minimum hate score of 51 points provided a good balance.
@@ 3.4.1 Simulated Annealing Algorithm
    Remove Irrelevant Candidates: Each candidate $C=\{{c}_{t}^{(i,j)}\}$ is then compared with each-other. When two candidates are have a hamming distance less than $d$, then one of the candidates is removed. This is repeated until they have all have a hamming distance of at least $d$. Furthermore, to avoid English answers, answers that have a high ratio of Latin characters to total characters are also removed to form the new set $\{\tilde{c}_{t}^{(i,j)}\}$
@@ 3.5 Human Annotation
Instructions For the main task, annotators were required to score each hate speech entry based on whether it qualifies as hate speech, counter-speech, or neither. If the sentence was determined to be hate speech, the annotator labeled it as ‘1’. If the sentence was counter-speech, it was labeled as ‘-1’. If the sentence did not fit into either category, it was labeled as ‘0’.
@@ 3.6 Analysis
Figure 3: The distribution of human labeling on hate-speech that has already been processed. This was generated from the first 800 instances of collected data.
Figure 4: A histogram showing the average ranking of human-preferred/written answers to AI generated answers. This was generated from the first 800 instances of collected data.
@@ 4.1 Creating the First HS–CS Pairs in Mandarin Chinese Using LLM-as-Judge
A principal goal was to harness an LLM-as-a-Judge (JudgeLM) to assist in producing paired HS–CS entries for Chinese. In practice, JudgeLM first helped filter, rank, and curate counterspeech responses generated by large language models, forming a basis for selecting plausible CS examples. This LLM-in-the-loop approach allowed us to rapidly develop a list of ~12,000 HS–CS pairs. Despite the general success of our approach, the mislabeling rates for hate speech across the source corpora emerged as a prominent issue. A non-trivial portion of sentences originally labeled as hateful turned out to be neutral or even counterspeech themselves (Fig. 3). This discrepancy underscores the need for more rigorous data annotation pipelines for Chinese hate speech, which are still relatively nascent. Moreover, in terms of the need for human-annotators, our pipeline was demonstrably not very cost effective; human annotators, in total, spend several hours processing and correcting AI generated responses, but were only able to create ~1000 out of the proposed 2,974 pairs of HS and CS. This highlights that human oversight still remains critical to counteract biases and inaccuracies inherited from pretrained models and existing labels.
@@ Appendix A Limitations
Another challenge arose from the use of a general-purpose language model, JudgeLM, tasked with rating the AI-generated counter-speech. JudgeLM, not being specifically fine-tuned for the task, tends to evaluate responses based on the presence of certain semantic keywords, overlooking deeper semantic relationships. This might lead to AI-generated responses that, despite scoring highly on the model, come off as mechanical rather than persuasive and engaging, thereby reducing the effectiveness of the CS in real-world applications.
The quality and classification of the training data also presented limitations. Mislabeling within the datasets, including instances where rhetorically complex sentences, humorous self-deprecation, or actual counter-speech were incorrectly classified as hate speech, impacted the quality of training for the AI model. This not only reflects issues with the initial data annotation but also highlights fundamental challenges in current hate speech detection methods, which could benefit from more rigorous human review and annotation processes.
@@
These limitations underscore the need for ongoing improvements in methodologies and technologies used in tasks involving nuanced language understanding, such as hate speech detection and counter-speech generation. Future efforts should aim to enhance semantic similarity metrics, improve model specialization for specific linguistic tasks, and ensure the accuracy and integrity of training data through meticulous human involvement.

### v2

@@ Abstract
Despite the global prevalence of Modern Standard Chinese language, counterspeech (CS) resources for Chinese remain virtually nonexistent. To address this gap in East Asian counterspeech research we introduce a corpus of Modern Standard Mandarin counterspeech that focuses on combating hate speech in Mainland China. This paper proposes a novel approach of generating CS by using an LLM-as-a-Judge, simulated annealing, LLMs zero-shot CN generation and a round-robin algorithm. This is followed by manual verification for quality and contextual relevance. This paper details the methodology for creating effective counterspeech in Chinese and other non-Eurocentric languages, including unique cultural patterns of which groups are maligned and linguistic patterns in what kinds of discourse markers are programmatically marked as hate speech (HS). In our analysis of the generated corpora, we provide strong evidence for the lack of open-source, properly labeled Chinese hate speech data and the limitations of using an LLM-as-Judge to score possible answers in Chinese. Moreover, the present corpus serves as the first East Asian language based CS corpus and provides an essential resource for future research on counterspeech generation and evaluation. [^1] The dataset can be found at github.com/
michaelbennieUFL/PANDA
Warning: The below text contains vulgar and oftentimes offensive speech. Any counterspeech or hate speech is used for exemplary purposes and doesn’t necessarily reflect the views of any researcher involved.
@@ 1 Introduction
Hate speech is typically characterized as any form of communication that demeans a specific group of people based on attributes like race, ethnicity, gender, sexual orientation, or religion (de Gibert et al., 2018). While HS may constitute a small proportion of social media content, its impact is significant, affecting nearly one-third of the population (Vidgen et al., 2019). The proliferation of hate speech on social media platforms has become a significant societal concern. While traditional approaches to mitigating HS have focused on content removal and moderation, these methods often raise concerns about freedom of speech. In response, counterspeech has emerged as a promising alternative strategy to combat HS while preserving free expression (Poudhar et al., 2024).
@@ 2.1 Hate Speech and Counterspeech
Counterspeech has gained traction as an alternative to content removal. Studies have demonstrated the efficacy of CS in enhancing online discourse quality and reducing the prevalence and impact of hateful behavior (Buerger, 2021). However, it’s important to note that the effectiveness of CS can vary significantly depending on the context and specific strategies employed. For example, the quantity of training data available to train an LLM on a specific language will predict the robustness of its generative function.
@@ 2.5 Current Evaluation Metrics
Datasets Open Source [^2] This paper used the Open Source Initiative’s definition of open source which can be found at opensource.org/osd. Total Instances HS/Offensive Speech Non-HS
@@ 3 Methodology
Figure 1: Proposed Data Processing Pipeline for Creating the Chinese Counterspeech Corpus. A1 through An refer to $n$ annotators that participated in this project.
@@ 3.1 Goals/Requirements
  • Creation of the First East Asian HS-CN Dataset. During our review of existing datasets, we identified significant gaps in Chinese counterspeech (CS) resources. Although datasets like Deng et al. (2022) and Zhou et al. (2022) include instances labeled as ’anti-bias’, their scope and definitions do not align with the specific focus of CS research. These datasets adopt a broader concept of ’anti-bias’, encompassing content that promotes fairness and addresses various forms of offensive language rather than specifically targeting hate speech. Our work addresses this gap by creating a dataset that exclusively targets hate speech and counter speech, providing a more focused resource for CS research.
@@
  • Freely Usable. All hate speech data collected for our dataset originate from open-source repositories. Additionally, we have released our model and the generated data under a permissive GPL license. This ensures that the generated and annotated data can be freely utilized in both commercial and non-commercial projects, promoting wider accessibility and application in various research and practical initiatives.
@@ 3.2 HS Sources
Three corpora ((Lu et al., 2023),(Zhou et al., 2022) and (Wang et al., 2022)) were later removed from the dataset due to restrictive licensing from them. What was left were 3 open-source datasets.
The COLDataset contains over 30,000 instances that are labeled either safe or offensive and, further, contains fine grained labels for each category Deng et al. (2022). The dataset was chosen due the fact that, under a cursory look, many, but not all, of the statements labeled offensive were in-fact hate speech. The second dataset used was ‘SexComment.csv’ from SWSR. This file focuses on finding and labeling sexist comments and also contains subcategories for the type of comment and whether it is targeted at an individual or a group Jiang et al. (2022). We decided to include this dataset to increase representation of sexist hate-speech in the database. The last dataset included was from CHSD which is actually a preprocessed dataset of HS that comes from COLD, CDIAL, and SWSR Rao et al. (2023).
@@ 3.3 Filtering of Data
The initial three corpora included entries that were labeled non-hate speech. In order to avoid the unnecessary computational cost of attempting to generate CS for non-HS sentences, we initially used some commands to filter out any rows that aren’t considered HS by the corpora. For CSHD, we removed any rows where ‘label’ equals ‘0.’ Likewise, for the COLD dataset, we kept only the data that had a label of ‘1’ and a sub label of ‘1’ (attacking individuals) or ‘2’ (attacking groups). For SWSR, we keep only the instances that had a label of ‘1’ (sexist) and all sub-categories except for ‘MA’ (micro aggressions) as we believed that it was harder to determine which answers counted as a hate-speech.
@@
We then use the scores given by the LLM and the text length to optimize over the set of possible subsets of hate-speech. As we wanted to have a subset that balances between high average hate score and a high average text length, we choose the metric of $log(AverageHateScore)*log(AverageTextLength)*NumInstances$. We limited the range from 500 to 3000 so that we would have a subset of answers that is large enough. As can be seen from Figure 2, we found that including strings that had a string length of at least 53 characters and a minimum hate score of 51 points provided a good balance.
@@ 3.4.1 Simulated Annealing Algorithm
    Remove Irrelevant Candidates: Each candidate $C=\{{c}_{t}^{(i,j)}\}$ is then compared with each-other. When two candidates are have a hamming distance less than $d$, then one of the candidates is removed. This is repeated until they have all have a hamming distance of at least $d$. Furthermore, to avoid English answers, responses that have a high ratio of Latin characters to total characters are also removed to form the new set $\{\tilde{c}_{t}^{(i,j)}\}$
@@ 3.5 Human Annotation
Instructions For the main task, annotators were required to score each hate speech entry based on whether it qualifies as hate speech, counterspeech, or neither. If the sentence was determined to be hate speech, the annotator labeled it as ‘1’. If the sentence was counterspeech, it was labeled as ‘-1’. If the sentence did not fit into either category, it was labeled as ‘0’.
@@ 3.6 Analysis
Figure 3: The distribution of human labeling on hate-speech that has already been processed. This was generated from the first 785 instances of collected data.
Figure 4: A histogram showing the ranking of human-preferred/written answers to AI generated answers. This was generated from the first 785 instances of collected data.
@@ 4.1 Creating the First HS–CS Pairs in Mandarin Chinese Using LLM-as-Judge
A principal goal was to harness an LLM-as-a-Judge (JudgeLM) to assist in producing paired HS–CS entries for Chinese. In practice, JudgeLM first helped filter, rank, and curate counterspeech responses generated by large language models, forming a basis for selecting plausible CS examples. This LLM-in-the-loop approach allowed us to rapidly develop a list of ~12,000 HS–CS pairs. Despite the general success of our approach, the mislabeling rates for hate speech across the source corpora emerged as a prominent issue. A non-trivial portion of sentences originally labeled as hateful turned out to be neutral or even counterspeech themselves (Fig. 3). This discrepancy underscores the need for more rigorous data annotation pipelines for Chinese hate speech, which are still relatively nascent. Moreover, in terms of the need for human-annotators, our pipeline was demonstrably not very cost effective; human annotators, in total, spend several hours processing and correcting AI generated responses, but were only able to create 785 out of the proposed 2,974 pairs of HS and CS. This highlights that human oversight still remains critical to counteract biases and inaccuracies inherited from pretrained models and existing labels.
@@ Appendix A Limitations
Another challenge arose from the use of a general-purpose language model, JudgeLM, tasked with rating the AI-generated counterspeech. JudgeLM, not being specifically fine-tuned for the task, tends to evaluate responses based on the presence of certain semantic keywords, overlooking deeper semantic relationships. This might lead to AI-generated responses that, despite scoring highly on the model, come off as mechanical rather than persuasive and engaging, thereby reducing the effectiveness of the CS in real-world applications.
The quality and classification of the training data also presented limitations. Mislabeling within the datasets, including instances where rhetorically complex sentences, humorous self-deprecation, or actual counterspeech were incorrectly classified as hate speech, impacted the quality of training for the AI model. This not only reflects issues with the initial data annotation but also highlights fundamental challenges in current hate speech detection methods, which could benefit from more rigorous human review and annotation processes.
@@
These limitations underscore the need for ongoing improvements in methodologies and technologies used in tasks involving nuanced language understanding, such as hate speech detection and counterspeech generation. Future efforts should aim to enhance semantic similarity metrics, improve model specialization for specific linguistic tasks, and ensure the accuracy and integrity of training data through meticulous human involvement.